# Effects of anisotropy in an anisotropic extension of $w$CDM model

### v1

@@ Abstract
In this paper, we derive observational constraints on an anisotropic $w$CDM model from observational data including Baryonic Acoustic Oscillations (BAOs), Cosmic Chronometer (CC), Big Bang Nucleosynthesis (BBN), Pantheon Plus (PP) compilation of Type Ia supernovae, and SH0ES Cepheid host distance anchors. We find that anisotropy is of the order $10^{-13}$, and its presence in the $w$CDM model reduces $H_{0}$ tension by $\sim 2\sigma$ and $\sim 1\sigma$ in the analyses with BAO+CC+BBN+PP and BAO+CC+BBN+PPSH0ES data combinations, respectively. In both analyses, the quintessence form of dark energy is favored at 95% CL.
@@ I Introduction
Our Universe is expanding with accelerated rate of expansion as observed from type Ia supernovae (SNe Ia) observations [1, 2]. Later on, various other observations such as the large scale structure (LSS), the Baryonic Acoustic Oscillations (BAOs) and the cosmic microwave background (CMB) supported this observation. A number of cosmological models have been proposed/investigated in the literature to help us comprehend the dynamics of the universe. But, among all cosmological models, $\Lambda$CDM has proven to be the simplest mathematical model which has widely been accepted by the cosmological community and referred to as the standard cosmological model of cosmology. Basically, this model is composed of two main parts: cold (non-relativistic) dark matter (CDM) which is the reason behind the structure formation, and dark energy (DE) in the form of cosmological constant ($\Lambda$) which causes the late time accelerated expansion of Universe. The $\Lambda$CDM model provides an excellent fit across a broad range of scales and epochs [3, 4, 5, 6, 7], and successfully describes late-time accelerated expansion of the Universe [8, 9]. Despite the excellent fit to the current available cosmological observations, the model faces several theoretical and observational challenges. For instance, the nature of DE is not known accurately so far, and within the $\Lambda$CDM paradigm, DE is regarded as the cosmological constant in its most basic form, lacking any solid physical foundation. Except for its typical gravitational interactions with other components, the exact nature of dark matter remains unknown. Also, there is no concrete explanation of ‘coincidence problem’ stating why, despite having very distinct cosmic evolutions, do the dark matter densities and the DE have the same order? Is this coincidence indicating the possibility for interaction between the dark sector components? Up to what extent the cosmological principle has been tested? Is the Universe homogeneous and isotropic at cosmic scales?
@@
Further, a few indications in the observations point towards the need to expand the $\Lambda$CDM model in order to account for the growing conflicts between measurements made at early (high redshifts) and late (low redshifts) Universe [10, 11]. The Hubble constant, $H_{0}$, which represents the Universe’s current rate of expansion, has the greatest statistically significant tension. The Hubble tension appears when we compare the value of $H_{0}$ predicted by CMB measurements within the $\Lambda$CDM framework and the direct local distance ladder measurements, that is, the one estimated by the Cepheid calibrated supernovae SNe Ia. In particular, the Hubble tension is referred to as the disagreement at $5\sigma$ between the latest SH0ES (Supernovae and $H_{0}$ for the equation of state (EoS) of DE) collaboration [12] constraint, $H_{0}^{\rm R22}=(73.04\pm 1.04){\rm\,km\,s^{-1}\,Mpc^{-1}}$ at 68% confidence level (CL), based on the supernovae calibrated by Cepheids, and the Planck collaboration [3] value, $H_{0}=\left(67.27\pm 0.60\right){\rm\,km\,s^{-1}\,Mpc^{-1}}$ at 68% CL. The Hubble tension may be defined as a difference between the observed value of $H_{0}$ from two sets of observations: (i) all of the direct late time $\Lambda$CDM independent measurements, and (ii) all of the indirect model dependent estimations at early times. In general, the value of $H_{0}$ obtained at late times is greater than the one obtained at early times.
@@
This disparity could indicate the existence of novel physics outside of the $\Lambda$CDM cosmology [13, 14, 15, 16, 17, 18]. In the literature, a number of extensions of $\Lambda$CDM have been suggested to resolve the $H_{0}$ tension. These extensions include but not limited to: DM-DE interactions [19, 20, 21, 22, 23, 24, 25]; decaying DM [26, 27, 28, 29]; introducing Early DE [30, 31, 32, 33, 34]; and introducing a sign-switching DE at intermediate redshifts ($z\sim 2$[35, 36, 37, 38, 39]. The current status on $H_{0}$ tension and possible solutions can be found in recent review articles [40, 41, 42, 43]
@@
The Cosmological Principle (CP), which asserts that the Universe is statistically homogenous and isotropic in space and matter on vast scales $(100\rm Mpc)$, is one of the fundamental tenets of $\Lambda$CDM cosmology. Mathematically, such a universe is described by the Friedmann-Lemaître-Robertson-Walker (FLRW) space-time metric, in which all three of the metric’s spatial orthogonal components are functions of cosmic time ($t$) exclusively. This is the primary space-time metric that facilitates the creation of the effective and remarkably predictive standard model of cosmology. However, observational data suggest that there are slight fluctuations in CMB intensities originating from various directions of sky  [44]. Gamma Ray Bursts (GRBs), Quasars, Galaxies, and SNe Ia are among the recent findings that suggest the Universe may be anisotropic and provide substantial observational evidence (see [45] and references therein). There are other interesting observational evidences, which have questioned the validity of CP, for instance, a strong evidence for a violation of the CP of isotropy is found by the authors in [46] after analysing the Planck Legacy temperature anisotropy data. This violation is likely to represent a statistical fluctuation of order $\sim 10^{-9}$. Also, see the recent and interesting review [47] where the authors emphasise differences and synergies that truly stimulate additional research in this field as they detail existing observational clues for departures from the predictions of CP. The spatially homogeneous and anisotropic Universe is expressed mathematically by homogeneous and anisotropic metrics and corresponding models are commonly known as Bianchi type models [48, 49]. The simplest of the various Bianchi type models is the Bianchi type I model. The empirical limitations on the Bianchi type I model might offer a platform for evaluating the precision of FLRW models characterising the current era of the universe. Various anisotropic generalizations of the standard $\Lambda$CDM model have been investigated in recent years. In [50], the authors analyze an anisotropic model with the CMB and BAO data by fixing the drag redshift as $z_{\rm d}=1059.6$ and the last scattering redshift as $z_{*}=1089.9$ and constrain the anisotropy parameter, $\Omega_{\rm\sigma 0}\lesssim 10^{-15}$. A similar investigation is done in [51] by considering anisotropic expansion and curvature together from different data sets and discovered very little present-day expansion anisotropy. Other anisotropic models are also investigated by using different observational data in different contexts, see [52, 53, 54, 55, 56, 57, 58]. In [59], a theoretical approach of spatially homogeneous universes with late-time anisotropy is discussed. Most recently, in [60] an anisotroic model, $\Lambda$CDM + $\Omega_{\rm\sigma 0}$ is investigated with the aim to derive CMB independent constraints on $H_{0}$ with BAO, BBN, CC, Pantheon Plus and SH0ES data. Following [60], we are motivated here to derive observational constraints on two simplest extensions of vanilla $\Lambda$CDM model: (i) an isotropic extension by taking a constant EoS of DE instead of cosmological constant, namely $w$CDM model; (ii) an anisotropic extension, namely $w$CDM + $\Omega_{\rm\sigma 0}$ by adding expansion anisotropy to $w$CDM model. It is well known that the nature of DE could be investigated through its EoS parameter. As stated above, the fundamental nature of DE is unknown and there is no strong physical basis for assuming DE in the form of cosmological constant whose density remains constant even in an expanding background. In this paper, we have assumed a constant EoS parameter of DE. Our primary goal in this work is to analyze the possible effects of anisotropy on the constant EoS parameter of DE and the Hubble constant $H_{0}$. For this, we derive observational constraints using the recent observational data and compare the results of $w$CDM and $w$CDM + $\Omega_{\rm\sigma 0}$ models with the results obtained in [60] for $\Lambda$CDM and $\Lambda$CDM + $\Omega_{\rm\sigma 0}$, respectively.
@@
The structure of the paper is as follows: In Section II, we describe the details of the considered models. In Section III, we provide a brief overview of the data sets and methodology used for the observational analysis of the models. In Section IV, we present the observational constraints on the models, and discuss the results. In Section V, we conclude with the main findings of our study.
@@ II GOVERNING EQUATIONS AND MODELS
where the left side of the above expression shows the Einstein tensor $G_{\mu\nu}$, the Ricci tensor $R_{\mu\nu}$, the Ricci scalar $R$, and the metric tensor $g_{\mu\nu}$. Additionally, the right side shows the Newton’s gravitational constant $G$ and the energy momentum tensor $T_{\mu\nu}$. Further, for a perfect fluid with energy density $\rho$ and pressure $p$, $T_{\mu\nu}$ takes the form
@@
In case of the perfect fluid matter distribution, it reduces to
@@
where the dot denotes its cosmic time($t$) derivative. The evolution of the energy density $\rho_{i}$ of a perfect fluid $i$ with pressure $p_{i}$, energy density $\rho_{i}$, and constant EoS $w_{i}=p_{i}/\rho_{i}$ is provided by the continuity equation (5), which is as follows:
@@
where $\rho_{i0}$ stands for the current value of $\rho_{i}$, that is, at $a=a_{0}=1$. Here, $a=a_{0}=1$ signifies the present day value of the cosmic scale factor $a$. From this point on, every quantity with the subscript 0 indicates its value in the present day universe. We consider that the universe is made up of the standard cosmic fluids, namely the radiation (photons and neutrinos) whose EoS is $w_{\rm r}=p_{r}/\rho_{r}=\frac{1}{3}$, pressureless fluid (baryonic and cold dark matter) whose EoS is $w_{\rm m}=p_{m}/\rho_{m}=0$, and DE fluid with a constant EoS $w_{\rm de}$, to be fixed by observations in the analysis. Further, assuming only gravitational interaction between these energy components, the continuity equation (5) being satisfied by each component separately, and in view of (6), this gives
@@
where $\Omega_{\sigma 0}=\frac{\sigma_{0}^{2}}{3H_{0}^{2}}$, $\Omega_{\rm r0}=\frac{8\pi G}{3H_{0}^{2}}\rho_{\rm r0}$, $\Omega_{\rm m0}=\frac{8\pi G}{3H_{0}^{2}}\rho_{\rm m0}$ and $\Omega_{\rm de0}=\frac{8\pi G}{3H_{0}^{2}}\rho_{\rm de0}$, respectively denote the expansion anisotropy, radiation, matter, and DE density parameters, satisfying $\Omega_{\rm\sigma 0}+\Omega_{\text{r}0}+\Omega_{\text{m}0}+\Omega_{\rm de0}=1$. Here, the expansion anisotropy parameter $\Omega_{\rm\sigma 0}$ is purely a geometric term which is non-negative.
@@
We follow standard neutrino scheme with normal hierarchy which has three species of neutrino, out of which two are massless and one is massive neutrino with the standard number of effective neutrino species, $N_{\rm eff}=3.046$ and minimum allowed mass, $m_{\nu}=0.06\,\rm eV$.
@@ III Data and methodology
where $H^{\text{obs}}(z_{i})$ signifies the value of the observed Hubble parameter, along with its corresponding standard deviation $\sigma^{2}_{H^{\text{obs}}(z_{i})}$ from the table provided. The equivalent $H^{\text{th}}(z_{i})$ represents the Hubble parameter value in theory obtained from the considered cosmological model.
@@
Pantheon Plus and SH0ES: Historically, Type Ia supernovae (SNe Ia) have played crucial role in developing the conventional model of the universe. These supernovae yield useful distance moduli measurements, which constrain the late-time expansion rate or the uncalibrated luminosity distance $H_{0}d_{L}(z)$. For a supernova at redshift $z$, the theoretical apparent magnitude $m_{B}$ is determined by the following equation:
@@
In the present study, we make use of SNe Ia distance modulus measurements from the Pantheon+ sample [78]. We name the collection as PP, which comprises 1701 light curves that correspond to 1550 unique supernovae Ia within the $z\in[0.001,2.26]$ redshift range. We also include constraints on $H_{0}$ and $M_{B}$ by including SH0ES Cepheid host distance anchors [78] into our analysis; this dataset is referred to as PPSH0ES.
We utilise PP data to minimise the equation that follows to constrain the cosmological parameters:
@@
While SNe analysis alone suffers from degeneracy between the parameters $M_{B}$ and $H_{0}$, we get over this restriction by adding the recently reported SH0ES Cepheid host distance anchors (R22) into the likelihood. With this changement, we are able to confine $M_{B}$ as well as $H_{0}$. The following adjustments are made to the SN distance residuals after including SH0ES Cepheid host distances:
@@
The baseline free parameters of the model $w$CDM+$\Omega_{\sigma 0}$ and $w$CDM are as discussed in section II. We have assumed three neutrino species, approximated as two massless states and a single massive neutrino of mass $m_{\nu}=0.06\,\rm eV$. We have used uniform priors: $\omega_{\rm b}\in[0.01,0.03]$, $\omega_{\rm c}\in[0.05,0.25]$, $\,H_{0}\in[60,80]$, $\Omega_{\sigma 0}\in[0,0.001]$, $w_{\rm de0}\in[-2,0]$ and the publicly available Boltzmann code CLASS [79] with the parameter inference code Monte Python [80] to obtain correlated Monte Carlo Markov Chain (MCMC) samples. We analyze the MCMC samples using the python package GetDist [^1] https://getdist.readthedocs.io/en/latest/intro.html.

### v2

@@ Abstract
In this paper, we derive observational constraints on an anisotropic $w$CDM model from datasets including Baryonic Acoustic Oscillations (BAOs), Cosmic Chronometer (CC), Big Bang Nucleosynthesis (BBN), Pantheon Plus (PP) compilation of Type Ia supernovae, and SH0ES Cepheid host distance anchors. We find that anisotropy is of the order $10^{-13}$, and its presence in the $w$CDM model reduces the $H_{0}$ tension by $\sim 2\sigma$ and $\sim 1\sigma$ in the analyses with BAO+CC+BBN+PP and BAO+CC+BBN+PPSH0ES data combinations, respectively. In both analyses, the quintessence form of dark energy is favored at 95% CL.
@@ I Introduction
Our Universe is expanding with an accelerated rate of expansion as observed from type Ia supernovae (SNe Ia) observations [1, 2]. Later on, various other measurements such as the large-scale structure (LSS), the Baryonic Acoustic Oscillations (BAOs), and the cosmic microwave background (CMB) supported this observation. Several theoretical frameworks have been proposed/investigated in the literature to help us comprehend the dynamics of the Universe. But, among all these, the $\Lambda$CDM model has proven to be the simplest mathematical framework widely accepted by the research community and referred to as the standard model of cosmology. It is composed of two major components: cold (non-relativistic) dark matter (CDM) which is the reason behind the structure formation, and dark energy (DE) in the form of cosmological constant ($\Lambda$) which causes the late time accelerated expansion of Universe. This $\Lambda$CDM paradigm provides an excellent fit across a broad range of scales and epochs [3, 4, 5, 6, 7], and successfully describes late-time accelerated expansion of the Universe [8, 9]. Despite the excellent fit to the current available cosmological datasets, the model faces several theoretical and observational challenges. For instance, the nature of DE is not known accurately so far. Within the standard $\Lambda$CDM scenario, DE is regarded as the cosmological constant in its most basic form, lacking any solid physical foundation. The exact nature of dark matter remains unknown except for its typical gravitational interactions with other components. Also, there is no concrete explanation of the ‘coincidence problem’ which states: why, despite having very distinct cosmic evolutions, do the dark matter and the DE densities have the same order at present? Is this coincidence indicating the possibility of interaction between the dark sector components? Up to what extent the cosmological principle has been tested? Is the Universe homogeneous and isotropic at cosmic scales?
@@
Further, a few indications in the observations point towards the need to expand the $\Lambda$CDM model to account for the growing conflicts between measurements made at early (high redshifts) and late (low redshifts) Universe [10, 11]. The Hubble constant, $H_{0}$, which represents the Universe’s current rate of expansion, has the greatest statistically significant tension. The Hubble tension appears when we compare the value of $H_{0}$ predicted by CMB measurements within the $\Lambda$CDM framework and the direct local distance ladder measurements, that is, the one estimated by the Cepheid calibrated SNe Ia. In particular, the Hubble tension is referred to as the disagreement at $5\sigma$ between the latest SH0ES (Supernovae and $H_{0}$ for the equation of state (EoS) of DE) collaboration [12] constraint, $H_{0}^{\rm R22}=(73.04\pm 1.04){\rm\,km\,s^{-1}\,Mpc^{-1}}$ at 68% confidence level (CL), based on the supernovae calibrated by Cepheids, and the Planck collaboration [3] value, $H_{0}=\left(67.27\pm 0.60\right){\rm\,km\,s^{-1}\,Mpc^{-1}}$ at 68% CL.
@@
This disparity could indicate the existence of novel physics outside of the $\Lambda$CDM cosmology [13, 14, 15, 16, 17, 18]. In the literature, many extensions of $\Lambda$CDM have been suggested to resolve the $H_{0}$ tension. These extensions include but not limited to: DM-DE interactions [19, 20, 21, 22, 23, 24, 25]; decaying DM [26, 27, 28, 29]; introducing Early DE [30, 31, 32, 33, 34]; and introducing a sign-switching DE at intermediate redshifts ($z\sim 2$[35, 36, 37, 38, 39]. The current status on $H_{0}$ tension and possible solutions can be found in recent review articles [40, 41, 42, 43]
@@
The Cosmological Principle (CP), which asserts that the Universe is statistically homogenous and isotropic in space and matter on vast scales $(\sim 100\rm Mpc)$, is one of the fundamental tenets of $\Lambda$CDM cosmology. Mathematically, such a Universe is described by the Friedmann-Lemaître-Robertson-Walker (FLRW) space-time metric, in which all three of the metric’s spatial orthogonal components are functions of cosmic time ($t$) exclusively. This is the primary space-time metric that facilitates the creation of the effective and remarkably predictive standard model of cosmology. However, observational data suggest that there are slight fluctuations in CMB intensities originating from various directions of sky  [44]. Gamma Ray Bursts (GRBs), Quasars, Galaxies, and SNe Ia are among the recent findings that suggest the Universe may be anisotropic and provide substantial observational evidence (see [45] and references therein). There are other interesting pieces of evidence, that have questioned the validity of CP, for instance, a piece of strong evidence for a violation of the CP of isotropy is found by the authors in [46] after analyzing the Planck Legacy temperature anisotropy data. This violation is likely to represent a statistical fluctuation of order $\sim 10^{-9}$. Also, see the recent and interesting review [47] where the authors emphasize differences and synergies that truly stimulate additional research in this field as they detail existing observational clues for departures from the predictions of CP. The spatially homogeneous and anisotropic Universe is expressed mathematically by homogeneous and anisotropic metrics and corresponding models are commonly known as Bianchi-type models [48, 49]. The simplest one among these is the Bianchi type I.
The empirical limitations on the Bianchi type I model might offer a platform for evaluating the precision of FLRW frameworks characterizing the current era of the Universe. Various anisotropic generalizations of the standard $\Lambda$CDM framework have been investigated in recent years. In [50], the authors analyze an anisotropic scenario with the CMB and BAO data by fixing the drag redshift as $z_{\rm d}=1059.6$ and the last scattering redshift as $z_{*}=1089.9$ and constrain the anisotropy parameter, $\Omega_{\rm\sigma 0}\lesssim 10^{-15}$. A similar investigation is done in [51] by considering anisotropic expansion and curvature together from different data sets and discovering little present-day expansion anisotropy. Other anisotropic models are also investigated by using different data sets in different contexts, see [52, 53, 54, 55, 56, 57, 58]. In [59], a theoretical approach to a spatially homogeneous Universe with late-time anisotropy is discussed. Most recently, in [60] an anisotropic extension, $\Lambda$CDM + $\Omega_{\rm\sigma 0}$ is investigated to derive CMB independent constraints on the Hubble constant with BAO, BBN, CC, Pantheon Plus and SH0ES data. Following [60], we are motivated here to derive constraints on two simplest extensions of vanilla $\Lambda$CDM paradigm: (i) an isotropic extension by taking a constant EoS of DE instead of a cosmological constant, namely the $w$CDM model; (ii) an anisotropic extension, namely $w$CDM + $\Omega_{\rm\sigma 0}$ by adding expansion anisotropy to the $w$CDM model. It is well known that the nature of DE could be investigated through its EoS parameter. As stated above, the fundamental nature of DE is unknown and there is no strong physical basis for assuming DE in the form of a cosmological constant whose density remains constant even in an expanding background. In this paper, we have assumed a constant EoS parameter of DE. Our primary goal in this work is to analyze the possible effects of anisotropy on the constant EoS parameter of DE and the Hubble constant $H_{0}$. For this, we derive the constraints on the free parameters of the considered models using recent data and compare our results with the results obtained in [60] for the $\Lambda$CDM and $\Lambda$CDM + $\Omega_{\rm\sigma 0}$, respectively.
@@
The structure of the paper is as follows: In Section II, we describe the mathematical details of the considered scenarios. In Section III, we provide a brief overview of the data sets and methodology used for the analysis of the models. In Section IV, we present the constraints on the parameters under study and discuss the results. In Section V, we conclude with the main findings of our study.
@@ II GOVERNING EQUATIONS AND MODELS
where the left side of the above expression shows the Einstein tensor $G_{\mu\nu}$, the Ricci tensor $R_{\mu\nu}$, the Ricci scalar $R$, and the metric tensor $g_{\mu\nu}$. Additionally, the right side shows Newton’s gravitational constant $G$ and the energy-momentum tensor $T_{\mu\nu}$. Further, for a perfect fluid with energy density $\rho$ and pressure $p$, $T_{\mu\nu}$ takes the form
@@
In the case of the perfect fluid matter distribution, it reduces to
@@
where the dot denotes its cosmic time($t$) derivative. The evolution of the energy density $\rho_{i}$ of a perfect fluid $i$ with pressure $p_{i}$ and constant EoS $w_{i}=p_{i}/\rho_{i}$ is provided by the continuity equation (5), which is as follows:
@@
where $\rho_{i0}$ stands for the current value of $\rho_{i}$, that is, at $a=a_{0}=1$. Here, $a=a_{0}=1$ signifies the present-day value of the cosmic scale factor $a$. From this point on, every quantity with the subscript 0 indicates its value in the present-day Universe. We consider that the Universe is made up of the standard cosmic fluids, namely the radiation (photons and neutrinos) whose EoS is $w_{\rm r}=p_{r}/\rho_{r}=\frac{1}{3}$, pressureless fluid (baryonic and cold dark matter) whose EoS is $w_{\rm m}=p_{m}/\rho_{m}=0$, and DE fluid with a constant EoS $w_{\rm de}$, to be fixed by observations in the analysis. Further, assuming only gravitational interaction between these energy components, the continuity equation (5) is satisfied by each component separately, and because of (6), this gives
@@
where $\Omega_{\sigma 0}=\frac{\sigma_{0}^{2}}{3H_{0}^{2}}$, $\Omega_{\rm r0}=\frac{8\pi G}{3H_{0}^{2}}\rho_{\rm r0}$, $\Omega_{\rm m0}=\frac{8\pi G}{3H_{0}^{2}}\rho_{\rm m0}$ and $\Omega_{\rm de0}=\frac{8\pi G}{3H_{0}^{2}}\rho_{\rm de0}$, respectively denote the expansion anisotropy, radiation, matter, and DE density parameters, satisfying $\Omega_{\rm\sigma 0}+\Omega_{\text{r}0}+\Omega_{\text{m}0}+\Omega_{\rm de0}=1$. Here, the expansion anisotropy parameter $\Omega_{\rm\sigma 0}$ is purely a geometric term that is non-negative.
@@
We follow the standard neutrino scheme with a normal hierarchy which has three species of neutrino, out of which two are massless and one is massive neutrino with the standard number of effective neutrino species, $N_{\rm eff}=3.046$ and minimum allowed mass, $m_{\nu}=0.06\,\rm eV$.
@@ III Data and methodology
where $H^{\text{obs}}(z_{i})$ signifies the value of the observed Hubble parameter, along with its corresponding standard deviation $\sigma^{2}_{H^{\text{obs}}(z_{i})}$ from the table provided. The equivalent $H^{\text{th}}(z_{i})$ represents the Hubble parameter value in theory obtained from the considered model.
@@
Pantheon Plus and SH0ES: Historically, Type Ia supernovae (SNe Ia) have played a crucial role in developing the conventional model of the Universe. These supernovae yield useful distance moduli measurements, constraining the late-time expansion rate or the uncalibrated luminosity distance $H_{0}d_{L}(z)$. For a supernova at redshift $z$, the theoretical apparent magnitude $m_{B}$ is determined by the following equation:
@@
In the present study, we use SNe Ia distance modulus measurements from the Pantheon+ sample [78]. We name the collection PP, which comprises 1701 light curves that correspond to 1550 unique supernovae Ia within the $z\in[0.01,2.26]$ redshift range. We also include constraints on $H_{0}$ and $M_{B}$ by including SH0ES Cepheid host distance anchors [78] into our analysis; this dataset is referred to as PPSH0ES.
We utilize PP data to minimize the equation that follows to constrain the cosmological parameters:
@@
While SNe analysis alone suffers from degeneracy between the parameters $M_{B}$ and $H_{0}$, we overcome this restriction by adding the recently reported SH0ES Cepheid host distance anchors (R22) into the likelihood. With this change, we can confine $M_{B}$ as well as $H_{0}$. The following adjustments are made to the SN distance residuals after including SH0ES Cepheid host distances:
@@
The baseline free parameters of the model $w$CDM+$\Omega_{\sigma 0}$ and $w$CDM are as discussed in section II. We have assumed three neutrino species, approximated as two massless states and a single massive neutrino of mass $m_{\nu}=0.06\,\rm eV$. We have used uniform priors: $\omega_{\rm b}\in[0.01,0.03]$, $\omega_{\rm c}\in[0.05,0.25]$, $\,H_{0}\in[60,80]$, $\Omega_{\sigma 0}\in[0,0.001]$, $w_{\rm de0}\in[-2,0]$ and the publicly available Boltzmann code CLASS [79] with the parameter inference code Monte Python [80] to obtain correlated Monte Carlo Markov Chain (MCMC) samples. We analyze the MCMC samples using the python package GetDist [^1] https://getdist.readthedocs.io/en/latest/intro.html. It is important to note here, we use the above-mentioned data in two combinations: BAO+CC+BBN+PP and BAO+CC+BBN+PPSH0ES, and abbreviated as $D_{1}$ and $D_{2}$, respectively throughout the forthcoming text.